OpenStudy (anonymous):

Among the 12 solar collectors on display at a trade show, 9 are flat-plate collectors and the others are concentrating collectors. If a person visiting the show randomly selects 4 of the solar collectors to check out, what is the probability that 3 of them will be flat plate collectors?

OpenStudy (anonymous):

Okay so for this problem I think I have to use the hypergeometric distribution formula: \[h(x;n,a,N)=\frac{ \left(\begin{matrix}a \\ x\end{matrix}\right) \left(\begin{matrix}N-a \\ n-x\end{matrix}\right)}{ \left(\begin{matrix}N \\ n\end{matrix}\right) }\] where N=12 (population size), n=4(sample size), a=9(number of successes of a flat plated collector) and the oribability x=3. When I plug the follow into the above formula I get h(3;4,9,12)=84*3/495=.5091 (rounded to four decimals). Is this correct?

OpenStudy (mayankdevnani):

correct

OpenStudy (anonymous):

alright thank you.

OpenStudy (mayankdevnani):

welcome